OpenStudy (anonymous):

Why do the goblins react with such great emotion when they see that Thorin has the sword they call Ocrist? A. To the goblins, the sword is a symbol of violence and terror. B. To the goblins, the sword is a symbol of victory and joy. C. To the goblins, the sword is a symbol of hope and love. D. To the goblins, the sword is a symbol of confusion and mistakes.

OpenStudy (anonymous):

\(\huge {\color{navy} {help~please?} }\)

OpenStudy (anonymous):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge\rlap{\qquad\qquad\mathbf{\;\,Great~ job!}}\bbox[5pt,gold,border:2px solid red]{\qquad\qquad\color{red}{\mathbf{Great~ job!}}\qquad\qquad\quad}\\ \Huge\rlap{\mathbf{\,\;Gold~ stars~ for~ you!}}\bbox[5pt,gold,border:2px solid red]{\color{red}{\mathbf{Gold~ stars~ for~ you!}}}\\ \Huge \begin{matrix} \color{gold}{★★★}\\ \color{gold}{★★}\\ \color{gold}{★★★} \end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge\rlap{\qquad\qquad\mathbf{\;\,Great~ job!}}\bbox[5pt,gold,border:2px solid red]{\qquad\qquad\color{red}{\mathbf{Great~ job!}}\qquad\qquad\quad}\\ \Huge\rlap{\mathbf{\,\;Gold~ stars~ for~ you!}}\bbox[5pt,gold,border:2px solid red]{\color{red}{\mathbf{Gold~ stars~ for~ you!}}}\\ \Huge \begin{matrix} \color{gold}{★★★}\\ \color{gold}{★★}\\ \color{gold}{★★★} \end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

its blackops2luvr my account is crashing

OpenStudy (anonymous):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge\rlap{\qquad\qquad\mathbf{\;\,solo!}}\bbox[5pt,gold,border:2px solid red]{\qquad\qquad\color{red}{\mathbf{solo!}}\qquad\qquad\quad}\\ \Huge\rlap{\mathbf{\,\;blackops2luvr!}}\bbox[5pt,gold,border:2px solid red]{\color{red}{\mathbf{blackops2luvr!}}}\\ \Huge \begin{matrix} \color{gold}{★★★}\\ \color{gold}{★★}\\ \color{gold}{★★★} \end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (anonymous):

.